OpenStudy (skullpatrol):

Prove: $$\frac{c}{1} = c$$

OpenStudy (anonymous):

Yo get back to mathstack exchange

OpenStudy (anonymous):

not sure if this valid but c/1 = (c*1)/1 = c * (1/1) = c * 1 = c

OpenStudy (anonymous):